OpenStudy (anonymous):

42.5 grams of an unknown substance is heated to 105.0 degrees Celsius and then placed into a calorimeter containing 110.0 grams of water at 24.2 degrees Celsius. If the final temperature reached in the calorimeter is 32.4 degrees Celsius, what is the specific heat of the unknown substance? Show or explain the work needed to solve this problem, and remember that the specific heat capacity of water is 4.18 J/(°C x g).

OpenStudy (anonymous):

@ganeshie8

OpenStudy (abmon98):

Use \[Q=m*c*\Delta(T)\] We heated Water using the unknown susbtance C=specific heat capacity of water (4.18 J g-1 °C-1), m = mass of water heated (kg), DeltaT = change in temperature of the water (°C) Q will represent the heat transferred from the 42.5 g of the unknown substance to Water.

OpenStudy (abmon98):

a rise in temperature is realized to Water but a decrease in temperature of unknown Substance.

OpenStudy (anonymous):

I dont understand

OpenStudy (abmon98):

When the heated 42.5 substance was placed heat energy was transferred to water and lead to a rise in the water's temperature. we calculate the quantity of heat transferred using the equation above

OpenStudy (anonymous):

How do i do the work

OpenStudy (abmon98):

plug in your givens mass of water=110 specific heat capcity of water is 4.18 and Delta T=(final Temp-Initial Temp)=(32.4-24.2)) q=mcDeltaT

OpenStudy (anonymous):

im still confused :/

OpenStudy (abmon98):

about?

OpenStudy (anonymous):

how to get everything and how to get it

OpenStudy (anonymous):

how to do it*

OpenStudy (abmon98):

When you heat water what happens to its temperature?

OpenStudy (anonymous):

it rises

OpenStudy (abmon98):

Great, in your question its the same thing a heated substance is placed in water. Before the heated susbtance was added water had an intial susbtance of 24.2. As the heated substance is placed temperature increases from 24.2 to 32.4

OpenStudy (abmon98):

initial temperature of 24.2 **

OpenStudy (anonymous):

so how will we write this all out as a equation

OpenStudy (abmon98):

Specific heat capacity the heat required to raise the temperature of the unit mass of a given substance by a given amount. The value is given in your question which is 4.18 J/(°C x g). Q=110*4.18*(32.4-24.2)

OpenStudy (anonymous):

so it would be 4.18 J/(°C x g). Q=110*4.18*(32.4-24.2)

OpenStudy (abmon98):

yes that's the quantity of heat transferred from the 42.5 grams of unknown substance to water

OpenStudy (anonymous):

So would that be the answer to the whole thing?

OpenStudy (abmon98):

no, what happened to the temperature of the unknown substance when it was placed in water?

OpenStudy (anonymous):

it rose

OpenStudy (abmon98):

the unknown substance lost some of its heat energy and was given to water so its temperature drops

OpenStudy (anonymous):

ok i still dont know what the answer is

OpenStudy (abmon98):

what was the final temperature recorded in calorimeter?

OpenStudy (anonymous):

32.4

OpenStudy (abmon98):

and what was the starting temperature of the 42.5 grams of unknown substance?

OpenStudy (anonymous):

105.0

OpenStudy (abmon98):

heat loss=heat gain so 42.5*C*(32.4-105)=110*4.18*(32.4-24.2)

OpenStudy (anonymous):

so thats the answer?

OpenStudy (abmon98):

-3085.5*C=3770.36 C=3770.36/-3085=1.22

OpenStudy (anonymous):

so how do we get that answer without explaining it?

OpenStudy (abmon98):

its 1.22

OpenStudy (anonymous):

-3085.5*C=3770.36 C=3770.36/-3085=1.22 so that is the whole process of getting that answer

OpenStudy (abmon98):

we divided the 3370.36(which represent the heat gained) no the part above too! without the above part of calculation you couldnt use heat lost=heat gained because you wouldnt have know the amount of heat transferred.

OpenStudy (anonymous):

so what is the whole calculation?

OpenStudy (abmon98):

Q=mc*Delta T Q=110*4.18*(32.4-24.2) Heat lost=Heat gained mc*DeltaT=3770.36 42.5*c*(32.4-105)=3770.36 solve for c

OpenStudy (anonymous):

c is 1.22

OpenStudy (abmon98):

yes

OpenStudy (anonymous):

Where did c come from? I thought is was this,: (4.18 J g-1 °C-1), so why would we have to solve for c?

OpenStudy (anonymous):

Wait a second....wouldnt C actually be 4.18 J/(° C × g)....